how long will a tank to be filled with water when three taps are opened at the same time?(one is used as a discharge) the two bringing in water take 6min and 4min respectively.the outlet takes 9 min.
Answer by sudhanshu_kmr(1152)


portion of tank will fill in one minute = 1/6 + 1/4 - 1/9 = 11/36
so, it will take 36/11 minutes to fill the tank completely.